# Probabilistic and Differentiable Wireless Simulation with Geometric Transformers

### v1

@@ D.2 Probabilistic modelling
Evaluation. We use the DDIM sampler using 100 timesteps for visualizations in Fig. 5 and for the error analysis in Fig. 8. To evaluate the variational lower bound in Tbl. 3, we follow [39] and evaluate $L_{vlb}:=L_{0}+L_{1}+\dots L_{T}$, where $L_{0}:=-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1})$, $L_{t-1}:=D_{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))$ and $L_{T}:=D_{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0}),p(\mathbf{x}_{t}))$. To be precise, for each sample $\mathbf{x}_{0}$ on the test set, we get a single sample $\mathbf{x}_{t}$ from $q$ and evaluate $L_{vlb}$ accordingly. Table 3 reports the mean of all $L_{vlb}$ evaluations over the test set.
Additional results. Fig. 8, shows the quality of samples from $p_{\theta}(h|F,t,r)$ as a function of the amount of available training data, where we average over $3$ samples for each conditioning input. It is worth noting that diffusion samples have a slightly higher error than the predictive models. This shows that the joint probabilistic modelling of the whole scene is a more challenging learning task than a deterministic forward model.
To further evaluate the quality of generated rooms, we analyze how often the model generates walls between the receiver and transmitter, compared to the ground truth. Precisely, we plot the distribution of received power versus the distance of transmitter and receiver in Fig. 9 and color each point according to a line of sight test. We can see that, overall, Wi-GATr has an intersection error of $0.26$, meaning that in $26\%$ of the generated geometries, line of sight was occluded, while the true geometry did not block line of sight between receiver and transmitter. This confirms that the diffusion model correctly correlates the received power and receiver/transmitter positions with physically plausible geometries. While an error of $26\%$ is non-negligible, we note that this task involves generating the whole geometry given only a single measurement of received power, making the problem heavily underspecified. Techniques such as compositional sampling [21] could overcome this limitation by allowing to condition on multiple receiver and received power measurements.

### v2

@@ E.2 Probabilistic modelling
Evaluation. We use the DDIM sampler using 100 timesteps for visualizations in Fig. 5 and for the error analysis in Fig. 9. To evaluate the variational lower bound in Tbl. 2, we follow Nichol & Dhariwal (2021) and evaluate $L_{vlb}:=L_{0}+L_{1}+\dots L_{T}$, where $L_{0}:=-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1})$, $L_{t-1}:=D_{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))$ and $L_{T}:=D_{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0}),p(\mathbf{x}_{t}))$. To be precise, for each sample $\mathbf{x}_{0}$ on the test set, we get a single sample $\mathbf{x}_{t}$ from $q$ and evaluate $L_{vlb}$ accordingly. Table 2 reports the mean of all $L_{vlb}$ evaluations over the test set.
Additional results. Fig. 9, shows the quality of samples from $p_{\theta}(h|{\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}})$ as a function of the amount of available training data, where we average over $3$ samples for each conditioning input. It is worth noting that diffusion samples have a slightly higher error than the predictive models. This shows that the joint probabilistic modelling of the whole scene is a more challenging learning task than a deterministic forward model.
To further evaluate the quality of generated rooms, we analyze how often the model generates walls between the receiver and transmitter, compared to the ground truth. Precisely, we plot the distribution of received power versus the distance of transmitter and receiver in Fig. 10 and color each point according to a line of sight test. We can see that, overall, Wi-GATr has an intersection error of $0.26$, meaning that in $26\%$ of the generated geometries, line of sight was occluded, while the true geometry did not block line of sight between receiver and transmitter. This confirms that the diffusion model correctly correlates the received power and receiver/transmitter positions with physically plausible geometries. While an error of $26\%$ is non-negligible, we note that this task involves generating the whole geometry given only a single measurement of received power, making the problem heavily underspecified. Techniques such as compositional sampling (Du et al., 2023) could overcome this limitation by allowing to condition on multiple receiver and received power measurements.